OpenStudy (anonymous):

1−sin2y/(1+cosy)+(1+cosy)/siny+siny/(1−cosy)

OpenStudy (anonymous):

1-1-cos^2y/cosy+1-cosy/siny-siny/1-cosy

OpenStudy (anonymous):

1-(1-cosy)(1+cosy)/cosy+1+cosy/siny+siny/1-cosy

OpenStudy (anonymous):

\[1-\sin^2y/(1+cosy)+(1+cosy)/siny+siny/(1-cosy)\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

1-(1-cosy)+(1+cosy)/siny+siny/(1-cosy)

ganeshie8 (ganeshie8):

question is not clear, could you pelase use latex..

OpenStudy (anonymous):

to find the value of the expression

OpenStudy (anonymous):

3rd one

ganeshie8 (ganeshie8):

\[\large \dfrac{1-\sin^2y}{1+\cos y}+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\] like this ?

OpenStudy (anonymous):

but 1- (sin^2y)/(1+cosy)

ganeshie8 (ganeshie8):

\[\large 1-\dfrac{\sin^2y}{1+\cos y}+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[1-(1-cosy)+(1+cosy)/siny+siny/(1-cosy)\]

ganeshie8 (ganeshie8):

\[\large 1-\dfrac{\sin^2y}{1+\cos y}+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\] \[\large 1-\dfrac{1-\cos^2y}{1+\cos y}+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\] \[\large 1-\dfrac{(1+\cos y)(1-\cos y)}{1+\cos y}+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\] \[\large \cos y+\dfrac{1+cosy}{\sin y}+\dfrac{\sin y}{1-\cos y}\]

ganeshie8 (ganeshie8):

fine, so far ?

ganeshie8 (ganeshie8):

what was the given answer ?